# Enhancing Table Recognition with Vision LLMs: A Benchmark and Neighbor-Guided Toolchain Reasoner

### v2

@@ 5.4. Tool Invocation Analysis
On the PubTabNet dataset, the NGTR framework exhibits a marked preference for the image upscaling tool. This preference is consistent with this dataset’s overall lower image quality characteristic, as improving image resolution can better support subsequent table recognition tasks. For the SciTSR and WTW datasets, the border enhancement tool is invoked more frequently. These two datasets primarily feature tables with complete borders, and enhancing border features enables VLLM to recognize the table structure information more accurately. Moreover, the detection and cropping tool has the highest invocation rate on the WTW dataset, which aligns with the dataset’s characteristics in the wild scenarios. In such images, tables often occupy only a portion of the frame or appear displaced, making this tool effective in locating and extracting the target areas. In summary, the experimental results demonstrate that the NGTR framework adaptively selects appropriate image processing tools based on the characteristics of different datasets, thereby fully showcasing its flexible planning capabilities in cross-dataset scenarios.
@@
Table 7. Tool invocation preference analysis. Values show the percentage of samples using each tool, with bold values marking the dataset where each tool is most frequently used.
@@
5.5. Hyperparameter Sensitivity Analysis
The NGTR framework contains two core parameters: the maximum length of the toolchain execution plan $L$ and the number of plans generated each time $N$. Based on GPT, Experimental results on the SciTSR dataset are presented in Figures 6.
Our experimental results indicate that a moderate toolchain length achieves an adequate balance between complexity and performance, as excessive toolchain length increases combinatorial complexity and limits processing performance, thereby affecting the framework’s ability to generate high-quality solutions. Similarly, generating a moderate number of execution plans effectively balances solution quality and generation efficiency, whereas generating too few or too many plans slightly reduces performance. Therefore, a moderate toolchain length and number of execution plans can balance complexity and performance well, providing valuable guidance for the tool invocation.
6. Conclusion and Limitation
This paper addressed the under-explored challenge of table recognition using VLLMs in a training-free reasoning paradigm. We proposed the NGTR framework, which enhanced input image quality through lightweight models and neighbor-guided tool invocation strategies. Extensive experiments demonstrated that NGTR significantly improved VLLM-based table recognition performance, addressing critical bottlenecks such as low-quality input images. This work not only established a benchmark for table recognition but also highlighted the potential of VLLMs in advancing table understanding, paving the way for future research and applications.
Despite the strengths of our framework, we acknowledge several limitations that warrant further investigation. Firstly, its performance depends on the underlying toolkit. Secondly, when the available set of neighbor candidates does not sufficiently cover a wide range of scenarios, selecting inappropriate neighbors may lead to suboptimal performance. Nevertheless, we believe the NGTR framework demonstrates strong generalizability, serving as a versatile approach for tool invocation for various domains.
@@ Appendix A Logical to Markup Sequence Conversion
We divide the entire conversion process into two stages. In the first stage, we perform a preliminary preprocessing of the logical location information, associating the logical positions with the corresponding cell content and storing them in a tabular data matrix. The detailed implementation is shown in Algorithm 1.
@@
In the second stage, we traverse the tabular data matrix row by row, gradually converting the stored logical information and cell content into a markup sequence, ultimately generating the output. The specific implementation is detailed in Algorithm 2.
@@ Appendix B Study of Table Size
In this section, we analyze the scale of the tables involved in the benchmark evaluations. We quantify the size of tables in the SciTSR and PubTabNet datasets by counting the number of cells in each table. Furthermore, this section presents experiments designed to evaluate the impact of table size on table recognition performance. GPT is selected as a representative of VLLMs for these experiments, with the distribution details and experimental results illustrated in Figure 7. In small scale tables, the model demonstrates stable performance on both the SciTSR and PubTabNet datasets. However, as the table size increased, the model’s performance experienced a moderate decline. This indicates that larger tables introduce longer context lengths, thereby affecting the VLLMs’ TR task performance.
Figure 7. Table size distribution and its impact on TR performance in SciTSR (left) and PubTabNet (right) Datasets
@@
Figure 8. A case study is conducted using an evaluation sample from PubTabNet
@@ Appendix C Case Study
Figure 9. A case study is conducted using an evaluation sample from SciTSR
@@ Appendix D Parameter Settings of VLLMs
We report the parameter settings of six VLLMs. For open-source VLLMs, we select Phi-3.5-Vision-Instruct [^2] https://azure.microsoft.com/en-us/products/phi/ (Phi) and Llama-3.2-90B-Vision-Instruct [^3] https://www.llama.com/ (Llama) for evaluation. For closed-source VLLMs, we evaluate GPT-4o-mini [^4] https://openai.com/index/gpt-4o-mini-advancing-cost-efficient-intelligence/ (GPT-mini), Qwen-VL-Max [^5] https://qwenlm.github.io/blog/qwen-vl/ (Qwen), GPT-4o [^6] https://openai.com/index/hello-gpt-4o/ (GPT) and Gemini-1.5-Pro  [^7] https://deepmind.google/technologies/gemini/pro/ (Gemini). Specifically, when generating multiple tool invocation plans, we set the temperature parameter to 0.8 to encourage the generation of more diverse tool invocation plans. For other experiments, the temperature parameter was kept at 0 to ensure the stability of the experimental results. Additionally, the top_p parameter was set to 0.2, and the n_samples parameter was set to 1.
@@ Appendix E Details of TEDS Metric
TEDS is designed to measure the similarity between two tree structures. Specifically, TEDS is used to calculate the similarity between the HTML tree of the real tag and the HTML tree of the predicted tag. When applying the TEDS metric, the HTML table format must first be converted into a tree structure. The similarity is then calculated using the following formula:
(6) $\vspace{-0.04in}\text{TEDS}(T_{a},T_{b})=1-\frac{\text{EditDist}(T_{a},T_{b})}{\max(|T_{a}|,|T_{b}|)},$
@@
Figure 10. Simplified prompt template for VLLMs on table recognition tasks.
Figure 11. Chain-of-Thought prompt template for VLLMs on table recognition tasks.
Appendix F Implementation Details of Benchmark
We present two prompt templates designed to guide VLLM in generating markup sequence outputs for TR tasks in Figures 10 and 11. Prompt 1 is a simplified template that facilitates the execution of TR tasks by VLLM. In contrast, Prompt 2 introduces the concept of Chain-of-Thought by explicitly planning the table recognition process for VLLM. In our experiments, the more complex Prompt 2 achieved a slight improvement in TR tasks performance. However, since our benchmark evaluation aims to reflect VLLM’s recognition capabilities more directly, we employ Prompt 1 as the prompt template in all experiments.
Figures 12-17 illustrate the prompt templates designed for our benchmark evaluation of hierarchical evaluation tasks. Each template corresponds to a specific hierarchical task, as detailed below:
@@ Appendix F Implementation Details of Benchmark
  • Row Index-based Data Recognition Task. (Figure 16)
  • Column Index-based Data Recognition Task. (Figure 17)
Figure 12. Prompt template for VLLMs on visual table size detection task.
Figure 13. Prompt template on merged cell detection task.
Figure 14. Prompt template for VLLMs on content-based cell recognition task.
Figure 15. Prompt template for VLLMs on index-based cell recognition task.
Figure 16. Prompt template for VLLMs on row index-based data recognition task.
Figure 17. Prompt template for VLLMs on column index-based data recognition task.
Figure 18. Prompt template for generating multiple tool invocation plans in the NGTR framework
Figure 19. Prompt template for the reflection-driven tool utilization module in the NGTR framework
Appendix G Implementation Details of NGTR
Figure 18 illustrates the prompt template utilized by our proposed NGTR framework for generating multiple tool invocation plans. By inputting the target image along with its tool descriptions and identifiers, we leverage the planning capabilities of VLLM to produce several tool invocation plans for subsequent modules. This module encompasses two adjustable hyperparameters: the maximum length $L$ of the toolchain execution plan and the number $N$ of plans generated per iteration.

### v3

@@ 5.4 Tool Invocation Analysis
We calculated the tool usage rates for each test dataset to investigate further the NGTR framework’s preferences in tool invocation (i.e., the proportion of samples that invoke a particular tool among all samples where tools are invoked). The results are presented in Table 6. On the PubTabNet dataset, the NGTR framework exhibits a marked preference for the image upscaling tool. This preference is consistent with this dataset’s overall lower image quality characteristic. The detection and cropping tool has the highest invocation rate on the WTW dataset, which aligns with the dataset’s characteristics in the wild scenarios. In such images, tables often occupy only a portion of the frame or appear displaced, making this tool effective in locating and extracting the target areas. In summary, the NGTR framework adaptively selects appropriate image processing tools based on the characteristics of different datasets, thereby fully showcasing its flexible planning capabilities in cross-dataset scenarios.
@@
Table 6: Tool invocation preference analysis. Values show the percentage of samples using each tool, with bold values marking the dataset where each tool is most frequently used.
@@
Figure 6: The impact of the number of tools (shown on the left) and the number of multi-paths (shown on the right) on performance.
5.5 Hyperparameter Sensitivity Analysis
The NGTR framework contains two core parameters: the maximum length of the toolchain execution plan $L$ and the number of plans generated each time $N$. As shown in Figure 6, a moderate toolchain length achieves an adequate balance between complexity and performance, as excessive toolchain length increases combinatorial complexity and limits processing performance, thereby affecting the framework’s ability to generate high-quality solutions. Similarly, generating a moderate number of execution plans effectively balances solution quality and generation efficiency, whereas generating too few or too many plans slightly reduces performance. Therefore, a moderate toolchain length and number of execution plans can balance complexity and performance well, providing valuable guidance for the tool invocation.
6 Conclusion and Limitation
This paper addressed the under-explored challenge of table recognition using VLLMs in a training-free paradigm. We proposed the NGTR framework, which enhanced input image quality through lightweight models and neighbor-guided tool invocation strategies. Extensive experiments demonstrated that NGTR significantly improved VLLM-based table recognition performance. This work not only established a benchmark for table recognition but also highlighted the potential of VLLMs in advancing table understanding, paving the way for future research and applications.
Limitation. Despite the strengths of our framework, we acknowledge several limitations that warrant further investigation. Firstly, its performance depends on the underlying toolkit. Secondly, when the available set of neighbor candidates does not sufficiently cover a wide range of scenarios, selecting an inappropriate neighbor may lead to suboptimal performance. Nevertheless, we believe the NGTR framework demonstrates strong generalizability, serving as a versatile approach for tool invocation for various domains.
Acknowledgements
This research was supported by grants from the grants of Provincial Natural Science Foundation of Anhui Province (No.2408085QF193), USTC Research Funds of the Double First-Class Initiative (No. YD2150002501), the National Key Research and Development Program of China (Grant No. 2024YFC3308200), the National Natural Science Foundation of China (62337001), the Key Technologies R & D Program of Anhui Province (No. 202423k09020039) and the Fundamental Research Funds for the Central Universities (No. WK2150110032).
@@ Appendix A Logical to Markup Sequence Conversion
We divide the entire conversion process into two stages. As shown in Algorithm 1, the first stage involves a preliminary preprocessing of the logical location information, associating the logical positions with the corresponding cell content and storing them in a tabular data matrix.
@@
In the second stage, we traverse the tabular data matrix row by row, gradually converting the stored logical information and cell content into a mark-up sequence, eventually generating the conversion result. The specific implementation is detailed in Algorithm 2.
@@ Appendix B Study of Table Size
In this section, we analyze the scale of the tables involved in the benchmark evaluations. We quantify the size of tables in the SciTSR and PubTabNet datasets by counting the number of cells in each table. Furthermore, this section presents experiments designed to evaluate the impact of table size on table recognition performance. GPT is selected as a representative of VLLMs for these experiments, with the distribution details and experimental results illustrated in Figure 7. In small scale tables, the model demonstrates stable performance on both the SciTSR and PubTabNet datasets. However, as the table size increased, the model’s performance experienced a moderate decline. This indicates that larger tables introduce longer context lengths, thereby affecting the VLLMs’ table recognition task performance.
Figure 7: Table size distribution and its impact on TR performance in SciTSR (left) and PubTabNet (right) Datasets
@@
Figure 8: A case study is conducted using a sample from PubTabNet.
@@ Appendix C Case Study
Figure 9: A case study is conducted using a sample from SciTSR.
@@ Appendix D Parameter Settings of VLLMs
We report the parameter settings of six VLLMs. For open-source VLLMs, we select Phi-3.5-Vision-Instruct [^7] https://azure.microsoft.com/en-us/products/phi/ (Phi) and Llama-3.2-90B-Vision-Instruct [^8] https://www.llama.com/ (Llama) for evaluation. For closed-source VLLMs, we evaluate GPT-4o-mini (GPT-mini), Qwen-VL-Max [^9] https://qwenlm.github.io/blog/qwen-vl/ (Qwen), GPT-4o [^10] https://openai.com/index/hello-gpt-4o/ (GPT) and Gemini-1.5-Pro  [^11] https://deepmind.google/technologies/gemini/pro/ (Gemini). Specifically, when generating multiple tool invocation plans, we set the temperature parameter to 0.8 to encourage the generation of more diverse tool invocation plans. For other experiments, the temperature parameter was kept at 0 to ensure the stability of the experimental results. Additionally, the top_p parameter was set to 0.2, and the n_samples parameter was set to 1.
@@ Appendix E Details of TEDS Metric
TEDS is designed to measure the similarity between two tree structures. Specifically, TEDS calculates the similarity between the HTML tree of the real tag and the HTML tree of the predicted tag. When applying the TEDS metric, the HTML table format must first be converted into a tree structure. The similarity is then calculated using the following formula:
$\text{TEDS}(T_{a},T_{b})=1-\frac{\text{EditDist}(T_{a},T_{b})}{\max(|T_{a}|,|T_{b}|)},$ (5)
@@
Appendix F Datasets
We selected three public datasets to evaluate our model’s performance across various table structures and conditions. These datasets provide a well-rounded basis for testing and comparing table recognition methods.
SciTSR is a dataset of the scientific literature containing 15,000 tables and their structural labels from LaTeX source files. It also has 12,000 training samples and 3,000 test samples. The images are high-quality and include complex tables.
PubTabNet is created by matching XML files with PDF files of scientific articles from the Open Access Subset of PubMed Central™. PubTabNet includes 500,777 training samples and 9,115 validation samples with their HTML representations. PubTabNet is for end-to-end model training, where the table area is turned into an image with a resolution of 72 pixels per inch (PPI). This low PPI reduces the computational load but lowers image quality.
WTW is a dataset of real-world scenes with 10,970 training samples and 3,611 test samples. It includes structural information for each table and extreme cases of real-world tables, such as tilt, overlap, blur, occlusion, and curvature. WTW focuses on recognizing structural information and does not provide labels for cell content.
Table 7: Characteristics of selected table recognition datasets.
Dataset Tables Source
Markup Sequence
Annotations
Textual
Annotations
Year
SciTSR 15K ArXiv $\times$ $\checkmark$ 2019
PubTabNet 568K PubMed $\checkmark$ $\checkmark$ 2020
WTW 14.6K Wild Scenarios $\times$ $\times$ 2021
Appendix G Implementation Details of Benchmark
We present two prompt templates designed to guide VLLM in generating markup sequence outputs for TR tasks in Figures 10 and 11. Prompt 1 is a simplified template that facilitates the execution of TR tasks by VLLM. In contrast, Prompt 2 introduces the concept of Chain-of-Thought by explicitly planning the table recognition process for VLLM. In our experiments, the more complex Prompt 2 achieved a slight improvement in the performance of table recognition tasks. However, since our benchmark evaluation aims to reflect VLLM’s recognition capabilities more directly, we employ Prompt 1 as the prompt template in all experiments.
Figure 10: Simplified prompt template for VLLMs on table recognition tasks.
Figure 11: Chain-of-Thought prompt template for VLLMs on table recognition tasks.
Figures 12-17 illustrate the prompt templates designed for our benchmark evaluation of hierarchical evaluation tasks. Each template corresponds to a specific hierarchical task:
@@ Appendix G Implementation Details of Benchmark
  • Index-based Row Recognition Task. (Figure 16)
  • Index-based Line Recognition Task. (Figure 17)
Figure 12: Prompt template on visual table size detection task.
Figure 13: Prompt template on merged cell detection task.
Figure 14: Prompt template on content-based cell recognition task.
Figure 15: Prompt template on index-based cell recognition task.
Figure 16: Prompt template for row index recognition task.
Figure 17: Prompt template for column index recognition task.
Figure 18: Prompt template for generating multiple tool invocation plans in the NGTR framework
Figure 19: Prompt template for the reflection-driven tool utilization module in the NGTR framework
Appendix H Implementation Details of NGTR
Figure 18 illustrates the prompt template utilized by our proposed NGTR framework for generating multiple tool invocation plans. By inputting the target image along with its tool descriptions and identifiers, we leverage the planning capabilities of VLLM to produce several tool invocation plans for subsequent modules. This module embeds two adjustable hyperparameters into the prompt template for process adjustment. The designed hyperparameters include the maximum length $L$ of the toolchain execution plan and the number $N$ of plans generated per iteration.